OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = 2. P(x) = 2x3 - 9x2 + 27 P(2) =

OpenStudy (anonymous):

P(2)=2.2.3-9.2.2+2.27 =12-36+54 =30

OpenStudy (anonymous):

thank you :D